OpenStudy (tortillabluechip):

3y + 5x = 11 I'm studying algebra to try to master it before the quarterly exam, can anyone help by explaining step by step?

OpenStudy (shadowlegendx):

Are we solving for a particular variable?

OpenStudy (shadowlegendx):

Putting it in slope intercept form?

OpenStudy (tortillabluechip):

We're trying to solve for x and y I believe. I just looked up algebra equations and wrote down a few like the ones I saw I was doing it class, there's no specific instructions.

OpenStudy (shadowlegendx):

If putting it in slope intercept form... Remember, slope intercept is... \[y = mx + b\] Lets begin \[3y + 5x = 11\] Lets start by isolating y, we do so by subtracting 5x from both sides We get, \[3y = -5x + 11\] We must further isolate y, so we divide the entire equation by 3. \[y = -\frac{ 5 }{ 3 }x + \frac{ 11 }{ 3 }\] We have solved for y

OpenStudy (skullpatrol):

Note: m is called the slope and b is called the y-intercept.

OpenStudy (mathmale):

I'd bet there were instructions somewhere for this problem. There are various mathematical operations that you can perform on 3y + 5x = 11; be sure you know which one is desired by whomever wrote this test or quiz.